OpenStudy (anonymous):

Is (-3, -1) a solution to the system of equations below? Show us how you arrived at your answer. 4x+y=-13 -2x-3y=-9

OpenStudy (anonymous):

second equation * 2 4x+ 6y = 18 minus 4x + y = -13 we get 5y = -5 ; y=-1 putting value of y in any of the equation we get x = -3 ;)

OpenStudy (anonymous):

where did you get the 4x+6y=18

OpenStudy (anonymous):

that is the second equation multiplied by -2

hartnn (hartnn):

to check whether (-3,-1) is the solution of the system, put x=-3 and y=-1 in BOTH the equations ans see whether left side equals right side. if it equals then the point is the solution, if it doesn't then the point is not the solution. Does this make sense ?

OpenStudy (anonymous):

-2x-3y=-9 multiplied by -2 gives 2x+6y=18

hartnn (hartnn):

....BOTH the equations *and see....

OpenStudy (anonymous):

gotcha?

OpenStudy (phi):

@miteshchvm you lost a minus sign when you solved these equations. notice that in the 2nd equation -2x-3y=-9 when we replace x with -3 and y with -1, we get -2* -3 + -3* -1= 6+3= 9 +9 does not equal -9, so (-3,-1) is not a solution to this system of equations.